OpenStudy (anonymous):

4x+4-7(x+1)=6x+2 solve for x

hartnn (hartnn):

distribute the 7 first can you tell me whats 7(x+1) = ..?

OpenStudy (anonymous):

-7x-1

hartnn (hartnn):

yeah, now we just combine like terms 4x+4 -7x-7 = 6x+2 try to solve further ?

OpenStudy (anonymous):

i did but it came out wrong. is it -3x+3=6x+2?

hartnn (hartnn):

4-7 is -3 actually :)

hartnn (hartnn):

-3x-3 = 6x+2

OpenStudy (anonymous):

ohhhh see my mistake now! little careless mistakes :p thanks. can you help me with a different problem?

hartnn (hartnn):

sure thing :)

OpenStudy (anonymous):

e^7x-1=(e^5)^x know it

hartnn (hartnn):

is the entire 7x-1 exponent ? or just 7x?

OpenStudy (anonymous):

the entire 7x-1

hartnn (hartnn):

so,